40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  PSL  Not logged in ELOG logo
Message ID: 536     Entry time: Fri Mar 4 22:28:39 2011
Author: frank, tara 
Type: DailyProgress 
Category: RefCav 
Subject: Q measurement for refcav's suspension (RCAV/ACAV) 

Today we measured Q for several modes of RCAV and ACAV's suspension the numbers are summarized below.

 We removed the insulation  foams on both cavities for Q measurement.

We bounce the beam on reference cavity's mirror and used the motion of the reflected beam to

observe the motion of the cavity. For RCAV, we used two PDs for shadow sensing on 2 directions (vertical and horizontal motions)

For ACAV, we use a QPD to observe the beam motion.  PSD from 4 signals are taken to identify the peaks

and compared with the beat signal.

rcav_motion.png

fig1,PSD from RCAV H,V motion

 

acav_motion.png

fig2, PSD from ACAV, H,V motion

 

compare.png

Fig3, beat measurement compared with PSD of ACAV and RCAV motion observed by reflected beams

on the mirrors' surfaces. The PSD is offset for comparison. We still have unknown peaks in the beat to investigate.

Most of the peaks we can excite by white noise on the table do not match with the beat signal

 

We used PZT attached on the table to excite each mode at its frequency and measured Q.

Beam movement can be observed by eyes for rough identification which suspension mode we excite

 RCAV

f                  Q              movement

3.15 Hz     150 -300        N/A

3.73           190           Vertical

4.3 Hz      220           Vertical

4.55 Hz     140          Horizontal

5.1            150                V

7               100               V

ACAV

f              Q        movement

4.3 Hz     120     V

3.4        N/A       H(?)*

 

For 3.4 Hz horizontal mode, it dies too fast even though we drove at the right frequency,

so Q cannot be measured properly. Its motion is also uncertain. We cannot see by eyes

but the signal on QPD shows that the motion is mostly horizontal at 3.4 Hz.

 

We will try to get the beat signal and excite each mode we found to check what modes

from which cavities have significant effects on beat signal. Then we will fix the motion of

that cavity. All insulation are back to normal (I misaligned the mirror behind RCAV when I

slide the back panel back, some alignment have to be done later.)

I'll check the temperature and make sure that we can

get both cavities locked for Monday.

 

Attachment 4: q21.png  15 kB  Uploaded Sat Mar 5 00:49:41 2011  | Hide | Hide all
q21.png
Attachment 5: q24.png  11 kB  Uploaded Sat Mar 5 00:49:49 2011  | Hide | Hide all
q24.png
Attachment 6: q27.png  12 kB  Uploaded Sat Mar 5 00:49:58 2011  | Hide | Hide all
q27.png
Attachment 7: q29.png  12 kB  Uploaded Sat Mar 5 00:50:18 2011  | Hide | Hide all
q29.png
Attachment 8: q31.png  10 kB  Uploaded Sat Mar 5 00:50:33 2011  | Hide | Hide all
q31.png
Attachment 9: q34.png  12 kB  Uploaded Sat Mar 5 00:50:51 2011  | Hide | Hide all
q34.png
ELOG V3.1.3-